OpenStudy (anonymous):

use f(x)=2x^2 to evaluate f(x+h)-f(x)/h

OpenStudy (anonymous):

lets take it step by step, although in a week you will be able to do this in your head

OpenStudy (anonymous):

k

OpenStudy (anonymous):

\(f(x)=2x^2\), so \(f(x+h)=2(x+h)^2\) and now some (careful) algebra

OpenStudy (anonymous):

\[f(x+h)=2(x+h)^2=2(x+h)(x+h)=2\left(x^2+2xh+h^2\right)\] \[=2x^2+4xh+2h^2\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now it is easy from here on in because \[f(x+h)-f(x)=2x^2+4xh+2h^2-2x^2=4xh+2h^2\] as \(2x^2-2x^2=0\)

OpenStudy (anonymous):

finally divide by \(h\) since each term has an \(h\) in it, they will cancel as follows \[\frac{f(x+h)-f(x)}{h}=\frac{4xh+2h^2}{h}=\frac{h(4x+2h)}{h}=4x+2h\]

OpenStudy (anonymous):

all details are there, hope it is clear

OpenStudy (anonymous):

yes! Thank you.